OpenStudy (anonymous):

PLEASE HELPP!!!! SHOW WORK 2√6 - 6√54

OpenStudy (anonymous):

Again, break down \[\sqrt{54}\] into \[(\sqrt{9})*(\sqrt{6})\] and take the square root of 9 which is 3 and multiply it by the 6 in front of it to give you\[(2\sqrt{6})-(18\sqrt{6})\]

OpenStudy (anonymous):

You need to find a way to factorize the thing. Try to find the factors of root(54) knowing that root(x)*root(y)=root(x*y)

OpenStudy (anonymous):

@Quantaliinuxite @DustyJ u guys r a great help but i just dont get this